write an equation of the sine function with the given amplitude, period, phase shift, and vertical shift.
Amplitude:3
period: 4pi
Phase shift: -1/2pi
vertical shift 1

Answer by lwsshak3(11628)

write an equation of the sine function with the given amplitude, period, phase shift, and vertical shift.
Amplitude:3
period: 4pi
Phase shift: -1/2pi
vertical shift 1
**
Equation for sin function:Asin(Bx-C), A=amplitude, Period=2π/B, Phase shift=C/B
For given problem:
A=3
period=4π=2π/B
B=2π/4π=1/2
Phase shift=-π/2=C/B
C=-Bπ/2=-(1/2)π/2=-π/4
Equation: y=3sin(x/2+π/4)+1